I need your help
Given f(x) = x^2 + 5 and g(x) = 5x – 2, solve the following functions:
15. f(6)
16. g(-2)

Found 2 solutions by stanbon, checkley71:
Answer by stanbon(48535)

Given f(x) = x^2 + 5 and g(x) = 5x – 2, solve the following functions:
15. f(6)= 6^2+5=36+5=41
16. g(-2)=5(-2)-2=-10-2=-12
cheers,
Stan H.

Answer by checkley71(8403)

F(X)=X^2+5
F(6)=6^2+5
F(6)=36+5
F(6)=41
------------------------------------
G(X)=5X-2
G(-2)=5*-2-2
G(-2)=-10-2
G(-2)=-12